OpenStudy (anonymous):

Find the definite integration of | (6-x)/(x^3) | dx from integral 1 to 7

hero (hero):

\[\int\limits_{1}^{7}\frac{6-x}{x^3}dx\] \[\int\limits_{1}^{7}-\frac{6-x}{x^3}dx\]

OpenStudy (anonymous):

oh then after would you add them?

hero (hero):

No, it's absolute value so you'll have two separate answers. You don't add them afterwards.

OpenStudy (anonymous):

because the x values are positive, would it just be the first one. because it only has one space for one answer to input

hartnn (hartnn):

ok, do u realize, the function (6-x)/(x^3) is positive from 1 to 6 and negative from 6 to 7 ?

hartnn (hartnn):

so u have to split your integral from 1 to 6 and 6 to 7 \(\int_1^6 ....dx + \int_6^7 ...dx \)

OpenStudy (anonymous):

ohhhh gotcha thanks i'll try it!

hartnn (hartnn):

remember to put a -*minus* sign for 2nd integral.

hero (hero):

I should have known it wasn't that easy